You have $390,000 invested in a well-diversified portfolio. You inherit a house that is presently worth $230,000. Consider the eturn old portfolio House 5t 151 18% 30% The correlation coefficient between your portfolio and the house is 0.3. a. What is the expected return and the standard deviation for your portfolio comprising your old portfolio and the house? (Do not round intermediate calculations. Round your final answers to 2 decimal places.) Expected return Standard deviation b. Suppose you decide to sell the house and use the proceeds of $230,000 to buy risk-free T-bills that promise a 11% rate of return. Calculate the expected return and the standard deviation for the resulting portfolio. (Hint Note that the correlation coefficient between any asset and the risk-free Tbills is zero) (Do not round intermediate calculations. Round your final answers to 2 decimal places.) Expected return Standard deviation
Answer #1

Solution:

a) The weights of investment are:

390000 390000+230000 Wold 0.63

230000 house 390000230000 0.37

The expected return is given below:

E(R_{p})=left ( w_{old} imes E(R_{old}) ight )+left ( w_{house} imes E(R_{house}) ight )

  (0.63 x 5%) + (0.37 x 15%)

3. 15% + 5.55%

=8.7%

And

SD(R_{p})=sqrt{w_{old}^2 sigma^{2}_{old}+w_{house}^2 sigma^{2}_{house}+2 imes w_{old} imes w_{house} imes ho imes sigma_{old} imes sigma_{house} }

v/0.632 x 0.182 + 0.372 x 0.32 + 2 x 0.63 x 0.37 x 0.30 x 0.18 x 0.3

0.18

b.

The weight is:

230000 = 0.37 Ur- 390000 + 230000

Now the expected return is:

E(R_{p})=left ( w_{r} imes E(R_{r}) ight )+left ( w_{h} imes E(R_{h}) ight )

  =left ( 0.37 imes 11% ight )+left ( 0.63 imes 15% ight )

  =13.52%

SD(R_{r})=sqrt{w_{r}^2 sigma^{2}_{r}+w_{h}^2 sigma^{2}_{h}}

=sqrt{0.37^2 imes 0.11^2+0.63^2 imes 0.15^2}

  = 0.10